1.12x + 1.08y = 0
x+y = 500
No clue where to start. Please help.

Answer by checkley71(8403)

X+Y=500 OR X=500-Y NOW SUBSTITUTE (500-Y) FOR X IN EQUATION #1
WE GET 1.12(500-Y)+1.08Y=0
560-1.12Y+1.08Y=0
-.04Y=-560
Y=-560/-.04
Y=14,000 THEN
X+14,000=500
X=500-14,000
X=-13,500
PROOF
1.12(-13,500)+1.08*14,000=0
-15,120+15,120=0
0=0